Anna is 12 years younger than elen if anna is 25 how old is elen translating verbal phrases to mathematical phrases
Found 2 solutions by math_helper, josgarithmetic:
Answer by math_helper(2461)


Let A be Anna's age
Let E be Elen's age
This part: "Anna is 12 years younger than Elen" means
A = E - 12
Given Elen's age, subtract 12 to get Anna's age. Keep in mind, Anna is the younger of the two, and this equation agrees with that.
Notice we can add 12 to both sides to get the equivalent
A + 12 = E
This allows you to find Elen's age if someone gives you Anna's age.
We can also write this as
E = A + 12
which looks a little nicer.
If A = 25 what do you think E is? Remember, you can use E = A + 12 ... just replace 'A' with 25...

Answer by josgarithmetic(39618)

----------------------------------------------------------
Anna is 12 years younger than elen if anna is 25 how old is elen
--------------------------------------------------------
--------------------------------------------------------------------
Anna is 12 years younger than Ellen. if Anna is 25, how old is Ellen?
-------------------------------------------------------------------

Make the substitution now, if you want.
25 is 12 less than E.
As an equation, 25=E-12, and solving this is simple basic arithmetic.